solve Absolute value of 2x-3 is less than or equal to 8x+1
Answer by lwsshak3(6490)

solve Absolute value of 2x-3 is less than or equal to 8x+1
**
|2x-3|≤8x+1
first condition
2x-3≤8x+1
-6x≤4
divide by -6 and switch inequality sign
x≥-4/6=-2/3
x≥-2/3 (reject, does not check)
second condition
-(2x-3)≤8x+1
divide by -1 and switch inequality sign
2x-3≥-8x-1
10x≥2
x≥2/10=1/5
x≥1/5
..
solution:
x≥1/5